What's the answer for :
3x + y = 1
x - 2y = 5

Answer by dabanfield(803)

1.) 3x + y = 1
2.) x - 2y = 5
From 1.) we know that y = 1 -3x.
Substituting 1-3x for y in 2.) we have:
x - 2*(1-3x) = 5
x - 2 + 6x = 5
7x = 7
x = 1
Substitute x = 1 in either 1.) or 2.) to determine the value of y.